OpenStudy (nikosis01):

What is the approximate circumference of a circle with a radius of 30 inches? Use π ≈ 3.14. A. 62.8 inches B. 94.2 inches C. 188.4 inches D. 47.1 inches

OpenStudy (dejmccray):

C=2*pi*r so C=3.14*2*30...2*30=60 so then multiply 60(3.14)=188.4 which is C

OpenStudy (anonymous):

C D. formula for circumference is : 2πr the radius is 30, and you are using 3.14 for pi so plug them in: 2(3.14)(30) do the multiplication: 60(3.14) 188.4

OpenStudy (nikosis01):

so c right

OpenStudy (anonymous):

DA ANSWER IS C

OpenStudy (nikosis01):

ok

OpenStudy (dejmccray):

yep C